Calculate the gross profit and cost of goods sold form the following information:

Net sales ₹ 9,00,000
Gross profit is 20% on cost
1
Expert's answer
2020-09-25T09:29:18-0400

Cost of goods sold=X

Net sales= Cost of goods sold + Gross Profit.

Thus, Gross Profit= 20/100(X)


₹ 9,000,000= X + 20/100X

₹ 9,000,000=6/5X

X=₹ 9,000,000(5/6)

X=₹ 7,500,000

Thus; Cost of goods sold = ₹ 7,500,000

Gross Profit= Net Sales- Cost of goods sold.

₹ 9,000,000- ₹ 7,500,000=₹ 1,500,000

Gross Profit=₹ 1,500,000